chavarria0831:

A student has a sample of isopropanol (C3H7OH) that has a mass of 78.6 g. The molar mass of isopropanol is 60.1 g/mol. How many moles of isopropanol are contained in this sample?

Nnesha:

\[\large\rm mass~of ~C_3H_7OH (g) \cdot \frac{1 mol ~of ~C_3H_7OH}{molar~mas ~of~ C_3H_7OH}\]

chavarria0831:

so whats the answer?

Nnesha:

\[\large\rm mass ~of~C_3H_7OH= 78.6 g ~~,~~ Molar~mass~of~C_3h_7OH=60.1 \frac{g}{mol}\] plugin these numbers into the expression i posted and then simplify =answer

chavarria0831:

oh okay thanks

Nnesha:

you're welcome. let me know what you get and if you're still confuse, feel free to speak up. : =))

Allison:

@Shadow

Shadow:

This question has been sufficiently answered so I will be closing it. If you have any further questions, please feel free to post here or open a new question.